OpenStudy (anonymous):

What is the solution to p over zero point three – 1 = 5? 20 13.3 1.2 1.8

OpenStudy (anonymous):

\[\huge \frac{z}{0.3}-1=5\]??

OpenStudy (anonymous):

ya thats right

OpenStudy (anonymous):

add 1 to both sides, then multiply both sides by \(.3\)

OpenStudy (anonymous):

so what is it

OpenStudy (anonymous):

what is five plus one?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

k good what is six times point 3 ?

OpenStudy (anonymous):

18

OpenStudy (anonymous):

not quite \[6\times 3=18\] but you need \(6\times .3\)

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

you don't get how to multiply \(6\) times \(0.3\) ?

OpenStudy (anonymous):

ohh is the answer 1.8

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

ohh ok thank you

OpenStudy (anonymous):

yw even blue girls get the cows

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

bye thanks